OpenStudy (anonymous):

2logaX-1/2logaY+5logaW Write the expression as a single logarithm.

OpenStudy (solomonzelman):

a is the base right?

OpenStudy (solomonzelman):

\[2\log_ax-1/2\log_ay+5\log_aw\]\[knowing~~~that,~~~loga+logb=\log(a \times b)~~~~and~~~~~loga-logb=\log(a \div b)\]\[and~~~b~\log(a)=\log(a^b)\] the second tip helps you apply the first tip.

OpenStudy (solomonzelman):

\[\log_a(x^2)-\log_a(y^{1/2})+\log_a(w^5)\] you used the second tip, now lets use the first tip(s) , the + - thingy, \[\log_a(x^2 \div y^{1/2} \times w^5)\]